# Natural Language Processing RELIES on Linguistics

### v4

@@ 5 Low-Resource Settings
The term “low-resource” has multiple meanings Hedderich et al. (2021). In NLP and ML, we wish to efficiently learn to generalize to new tasks, domains, and languages, including global and historic languages, where feedback and data is sparse or nonexistent.
@@
Linguistics is crucial in situations where the amount of language data is limited, e.g., when developing technologies for languages with a limited amount of available recordings or written data. We can improve generalization performance across an array of LLM models by exploiting linguistic features in training (Zhang et al., 2024) or dictionaries for balanced data augmentation (Lu et al., 2024); also with benefits to speech processing (Kim, Jeon, and Lee, 2024). Processing particular types of language/particular languages can also pose challenges that require input from linguistics. For examples, polysynthetic languages such as those spoken in Canada (Gupta and Boulianne, 2020a, b) pose problems for automatic speech recognition (ASR): if ‘word’ is taken as the basic unit for recognition (which makes sense for most languages, but not these ones) the out-of-vocabulary (OOV) rate on new data is extreme.
At first glance, Tanzer et al.’s (2024) study may appear as a potential ‘linguistics-free’ avenue for addressing some extremely low resource-scenarios, since it suggests that LLMs could, to a certain extent, translate a language from one linguist’s grammar book instead of a corpus. But by contrary, the study shows crucial reliance on linguistics: a linguist is required to assess LLM performance, and even the largest LLMs considerably underperform relative to a linguist (unfamiliar with the language) who has read the book. Of course, it is also the case that the fundamental resource was created by a linguist.
Even when more data is available, linguistics can be of value: It can inform data selection in multi-lingual pretraining Ogunremi, Jurafsky, and Manning (2023), or more language-balanced tokenization Creutz and Lagus (2002); Limisiewicz et al. (2024); Fusco et al. (2024). [^8] Some research also suggests that “simply scaling up the number of languages […] in the pretraining is unhelpful” for generalization (Adelani et al., 2024), and “bigger is not always better” Wilcox et al. (2024).
@@
According to some predictions, half of the world’s spoken languages might be on the brink of extinction (UNESCO, 2023). Thus, there is scope for NLP tooling to document and revitalize endangered [^9] Terms like ‘endangered’ and ‘documentation’ are widespread in the literature we are surveying, though we also wish to point at a critique of this terminology, suggesting ‘reclamation’ instead (Leonard, 2020). languages (e.g., van Esch, Foley, and San, 2019; Levow, Ahn, and Bender, 2021; San et al., 2022), ideally through collaborations involving community members, linguists, and NLP practitioners. Linguistic expertise is relevant for structurally collecting and documenting language data, as well as helping to distill knowledge about the language into dictionaries and rules. Recent work has brought community members and NLP practitioners together for the purpose of language documentation through provisioning access to speech tools such as ASR (San et al., 2022), building datasets on online platforms (Everson, Honore, and Grimm, 2019; Zuckermann et al., 2021), and analysis of phonetic data (Kempton, 2017).
@@
Systems that require less compute, particularly at inference time, are of great interest. Critically, off-the-shelf conventional parsers/taggers defend their status as the go-to method in some situations, like for a certain forms of data analysis (e.g., counting the most common nouns), or affordable processing of huge corpora (e.g., from sentence segmentation to OpenIE (Angeli, Johnson Premkumar, and Manning, 2015)).
Effective engineering solutions to efficiency are achieved through quantization Dettmers et al. (2022) or low-rank adaptation Hu et al. (2022) of LLMs. However, soft inductive linguistic biases might be complementary, or an alternative in specific situations. Tailored evaluation metrics could help assess any incurred tradeoffs Zhou et al. (2022).
@@
Developers of technologies for under-resourced languages do not only encounter data scarcity, but probably also scarcity of input from native speakers who could help oversee a system’s design and application. Indeed, releasing technology in a top-down fashion can be harmful to local communities Bird (2022); Doğruöz and Sitaram (2022); Bird and Yibarbuk (2024). Thus, when applied with care, (field-)linguistics can help become aware of the diverse cultural contexts of local language communication situations, resulting in more sensitive NLP technologies, or the abstention from releasing harmful ones (e.g., when there is no community oversight/feedback).
@@ 6 Interpretability and Explanation
As the field concerned with systems of human language, linguistics offers a metalanguage—in the form of jargon and formalisms—for precise characterization of NL phenomena. The field of NLP draws on this metalanguage, as can be seen in the literature in observations like “phonemic representations exhibit higher similarities between languages compared to orthographic representations” (Jung et al., 2024) and “difficult to score [in MT evaluation] are the transitive past progressive, multiple connectors, and the ditransitive simple future I for English to German, and pseudogapping, contact clauses, and cleft sentences for English to Russian” (Avramidis et al., 2024).
@@
For understanding computational models of language such as LLMs with billions and trillions of parameters, directed binding methods are being developed that learn to relate model internals and processes to specific human-understandable descriptions, that often lie within our vocabulary of linguistic metalanguage. We will consider a few examples of such methods that particularly use a linguistic notion of binding: Rassin et al. (2023) bind linguistic concepts in V&L diffusion models with syntax trees and attention maps; Di Marco, Hämmerl, and Fraser (2023) and Jumelet and Zuidema (2023) detect linguistic feature representations in LLMs with prompts; Opitz and Frank (2022) bind parts of neural embedding representations to semantic aspects; Geva et al. (2022) view model decisions as constructed from concepts in a vocabulary. Explanations and binding can also adopt an outside, behavioral view Beguš, Dąbkowski, and Rhodes (2023); Behzad et al. (2023); Chang and Bergen (2024). For instance, Muñoz-Ortiz, Gómez-Rodríguez, and Vilares (2023) contrast linguistic patterns in human- and LLM-generated text in morphological, syntactic, and sociolinguistic aspects. Since such methods can be also seen as diagnostic methods, the section on Evaluation (section 4) is also relevant.
@@
On a bigger picture, linguistic metalanguage can help us understand, or establish hypotheses, for what happens in complex NLP systems, and how they differ. An example is the hypothesis that a pre-trained LM can be viewed through the lens of a classical NLP pipeline, which has received arguments for and against (Tenney, Das, and Pavlick, 2019; de Vries, van Cranenburgh, and Nissim, 2020; Niu, Lu, and Penn, 2022). The learning trajectory of LLMs can be traced linguistically, e.g., noting the point of syntax acquisition (Chen et al., 2024). LLMs have also been interpreted as “models of varieties of language” (a sociolinguistic notion, Grieve et al., 2025).
@@
Building a bridge to Resources (§3), we can use data sets that elicit linguistic phenomena to “open the black box” of LLMs down to the level of single neurons, as exemplified by Niu et al. (2024), who study neuron-level knowledge with the BLiMP corpus of linguistic minimal pairs (Warstadt et al., 2020). With similar means, Wang, Hu, and Zhao (2024) show that BERT-like models focus on sentence-level features, whereas LLMs such as GPT or Llama are sensitive to conventions, language complexity, and organization. Experiments with complex and simple language structures can illuminate learning environments for LLMs (Qin, Saphra, and Alvarez-Melis, 2024).
@@
Linguistically-agnostic interpretability methods can be mechanistic, i.e., formalizing the algorithms learned by neural networks Weiss, Goldberg, and Yahav (2021), or based on attributing a prediction to inputs via integration Sundararajan, Taly, and Yan (2017), classifiers Ribeiro, Singh, and Guestrin (2016), or Shapley values Shapley (1951). However, still, for interpreting and evaluating the results of such methods, we rely on linguistic analysis Schuff et al. (2022); Feldhus et al. (2023); Moeller, Nikolaev, and Padó (2024), to verify that an interpretability method “outputs meaningful explanations” Mardaoui and Garreau (2021) or “emphasize[s] specific types of linguistic compositions” (Kobayashi et al., 2024). This means that linguistic analyses are used to validate interpretability methods whose algorithms do not explicitly incorporate such representations.
@@
Finally, at the thousand-foot view, we see ideas from linguistics and adjacent fields take center stage in debates about how to interpret what NLP models are capable of representing—like how to define machine ‘understanding’ (Dunietz et al., 2020; Ray Choudhury, Rogers, and Augenstein, 2022) and if grounding is required for a model to capture meaning (e.g., Bender and Koller, 2020; Merrill et al., 2021; Pavlick, 2023).
@@ 7 Study of Language
Our last facet highlights linguistics (and related areas) as the application domain. That is, those who study language, but are not necessarily computational linguists themselves, are a user base that can motivate NLP tasks and tools. [^10] Unlike the previous categories, which highlight ways that linguistics contributes to NLP, we view this category as bidirectional: language study motivates research towards developing NLP tools which can then contribute to the study of language. To be clear, we regard the entirety of cL—spanning formal, descriptive, distributional/statistical, and experimental inquiries—as part of linguistics. (This includes the development and use of frameworks and resources of the sorts discussed in section 3.) These inquiries contribute to linguistic understanding by advancing accounts of linguistic diversity and the precision and robustness of theories. Naturally, NLP models and representations are part of the methodology of this research (in particular, the potential relevance of LLMs to broader linguistics and cognitive science has been widely debated: see Futrell and Mahowald (2025) for a review). For purposes of this section, though, we focus on specialized NLP software that can be applied to language studies, not NLP models and representations generally. These language-system-focused applications can be distinguished from most contemporary NLP applications, which seek to make content more accessible through language (for example, by translating it, summarizing it, or reasoning about it).
@@
Classic cL tasks like parsing have applications to corpus linguistics: in order to study a linguistic pattern, it is often important to query not just strings, but with grammatical abstractions like tags, dependencies, and phrases. This was not the original goal of parsing researchers: parsing was a cL topic before the days of substantial digital corpora, and was studied to examine computational capacity to model natural language grammar, highlighting issues like formal expressive power (Steedman, 2011). Now, the tools produced have found a secondary application in corpus linguistics, with parser outputs accessible through corpus search engines (Resnik and Elkiss, 2005; Ghodke and Bird, 2012; Bender et al., 2012; Hundt, Denison, and Schneider, 2012; Krause and Zeldes, 2016; Guillaume, 2021; Kulick, Ryant, and Santorini, 2022). [^11] Search in parsed corpora may also be motivated by information extraction applications (Shlain et al., 2020). The application to corpus linguistics motivates new angles on parsing research (e.g., syntactically searching a large text collection without having to pre-parse it in its entirety, or in a way that accounts for parser uncertainty).
@@
Documentary and historical linguistics also motivate and contribute to NLP advancement. In documentary and historical linguistics, data may be sparse, fragmented, primarily in image or audio form (without transcriptions), and lacking a standard orthography; the basic grammar of the language may still be a mystery; and the language may have few or no living speakers. (The case of endangered language documentation was discussed above in section 5.) The exigencies of such settings call for considerably more noise-tolerance and interactivity than run-of-the-mill NLP tasks. Pure-text tasks like normalization (Robertson and Goldwater, 2018), as well as signal processing capabilities like OCR for historical documents (Berg-Kirkpatrick, Durrett, and Klein, 2013) or low-resource speech processing (Duong et al., 2016; Anastasopoulos et al., 2017; Liu, Spence, and Prud’hommeaux, 2022), are important here. Algorithms and software infrastructure to assist humans engaging in multiple tasks collaboratively (elicitation, transcription, annotation, discovering grammatical generalizations, and developing usable resources for scholars and community members) in limited-data multimodal conditions are needed, and there is a long way to go before they can be used seamlessly by non-computational linguists (Gessler, 2022; Moeller and Arppe, 2024).
@@
Language teaching, as studied in the field of Applied Linguistics and practiced in classrooms as well as other modes of instruction and technological support, is another area ripe for further engagement with NLP. Much of the research in the NLP community has been formulated in narrow tasks like grammatical error correction (as a generation task) or essay scoring (as a regression task) (Burstein and Chodorow, 1999; Massung and Zhai, 2016; Wang et al., 2021; Klebanov and Madnani, 2022). Applications with richer contexts—for example, technologies that would monitor student progress over time and deliver adaptive pedagogical experiences—are perhaps more difficult to study, but could enhance student learning more holistically (Cui and Sachan, 2023; Qian et al., 2023; Glandorf et al., 2025; Bang et al., 2024).
@@
Finally, we note the potential for NLP to embrace applications facilitating language-focused study in fields beyond linguistics proper, including literature Jakobson (1987), education, law Marmor (2014), communications Miller (1951), translation studies Gambier and Van Doorslaer (2009), history Piotrowski (2012), argumentation Cohen (1987), and lexicography McArthur (1998). In the NLP community, these fields are associated and developed with active workshops, e.g., LaTeCH-CLfL (CL for Cultural Heritage, Social Sciences, Humanities and Literature—8th iteration in 2024) (Bizzoni et al., 2024); BEA (Building Educational Applications—19th iteration) Kochmar et al. (2024); Argument Mining (11th iteration) Ajjour et al. (2024); and NLLP (Natural Legal Language Processing–3rd iteration) Aletras et al. (2024), among others. Key objectives featured in these workshops include creating structured resources, increasing interpretability, evaluating models and resources, and coping with low-resource settings. This underscores the importance of interdisciplinary work between all of NLP, linguistics, and other fields from humanities and social sciences.

### v5

@@ 5 Low-Resource Settings
In NLP and ML, we aim to efficiently learn to generalize to new tasks, domains, and languages, including global and historic languages, where feedback and data is sparse or nonexistent (hedderich-etal-2021-survey).
@@
Linguistics is crucial in situations where the amount of language data is limited, e.g., when developing technologies for languages with a limited amount of available recordings or written data. We can improve generalization performance across an array of LLM models by exploiting linguistic features in training (zhang-etal-2024-hire) or dictionaries for balanced data augmentation (lu-etal-2024-llamax); also with benefits to speech processing (kim-etal-2024-audio). Processing particular types of language/particular languages can also pose challenges that require input from linguistics. For examples, polysynthetic languages such as those spoken in Canada (gupta-boulianne-2020-automatic; gupta-boulianne-2020-speech) pose problems for automatic speech recognition (ASR): if ‘word’ is taken as the basic unit for recognition (which makes sense for most languages, but not these ones) the out-of-vocabulary (OOV) rate on new data is extreme.
At first glance, tanzer2024a’s (tanzer2024a) study may appear as a potential ‘linguistics-free’ avenue for addressing some extremely low resource-scenarios, since it suggests that LLMs could, to a certain extent, translate a language from one linguist’s grammar book instead of a corpus. However, the study shows crucial reliance on linguistics: a linguist is required to assess LLM performance, and even the largest LLMs considerably underperform relative to a linguist (unfamiliar with the language) who has read the book. Of course, it is also the case that the fundamental resource was created by a linguist.
Even when more data is available, linguistics can be of value: It can inform data selection in multi-lingual pretraining ogunremi-etal-2023-mini, or more language-balanced tokenization creutz-lagus-2002-unsupervised; limisiewicz2024myte; fusco2024recurrent. [^8] Some research also suggests that “simply scaling up the number of languages […] in the pretraining is unhelpful” for generalization (Adelani et al., 2024), and “bigger is not always better” wilcoxbigger.
@@
According to some predictions, half of the world’s spoken languages might be on the brink of extinction (UNESCO.org_2023). Thus, there is scope for NLP tooling to document and revitalize endangered [^9] Terms like ‘endangered’ and ‘documentation’ are widespread in the literature we are surveying, though we also wish to point at a critique of this terminology, suggesting ‘reclamation’ instead (Leonard:20). languages (e.g., van-esch-etal-2019-future; levow-etal-2021-developing; san-etal-2022-automated), ideally through collaborations involving community members, linguists, and NLP practitioners. Linguistic expertise is relevant for structurally collecting and documenting language data, as well as helping to distill knowledge about the language into dictionaries and rules. Recent work has brought community members and NLP practitioners together for the purpose of language documentation through provisioning access to speech tools such as ASR (san-etal-2022-automated), building datasets on online platforms (everson-etal-2019-online; zuckermann-etal-2021-lara), and analysis of phonetic data (kempton-2017-cross).
@@
Systems that require less compute, particularly at inference time, are of great interest. Critically, off-the-shelf conventional parsers/taggers defend their status as the go-to method in some situations, like for a certain forms of data analysis (e.g., counting the most common nouns), or affordable processing of huge corpora (e.g., from sentence segmentation to OpenIE (angeli-etal-2015-leveraging)).
Effective engineering solutions to efficiency are achieved through quantization NEURIPS2022_c3ba4962_quant or low-rank adaptation hu2022lora of LLMs. However, soft inductive linguistic biases might be complementary, or an alternative in specific situations. Tailored evaluation metrics could help assess any incurred tradeoffs zhou-etal-2022-assessing.
@@
Developers of technologies for under-resourced languages do not only encounter data scarcity, but probably also scarcity of input from native speakers who could help oversee a system’s design and application. Indeed, releasing technology in a top-down fashion can be harmful to local communities bird-2022-local; dogruoz-sitaram-2022-language; bird-yibarbuk-2024-centering. Thus, when applied with care, (field) linguistics can help become aware of the diverse cultural contexts of local language communication situations, resulting in more sensitive NLP technologies, or the abstention from releasing harmful ones (e.g., when there is no community oversight/feedback).
@@ 6 Interpretability and Explanation
As the field concerned with systems of human language, linguistics offers a metalanguage—in the form of jargon and formalisms—for precise characterization of linguistic phenomena. The field of NLP draws on this metalanguage, as can be seen in the literature in observations like “phonemic representations exhibit higher similarities between languages compared to orthographic representations” (jung-etal-2024-mitigating) and “difficult to score [in MT evaluation] are the transitive past progressive, multiple connectors, and the ditransitive simple future I for English to German, and pseudogapping, contact clauses, and cleft sentences for English to Russian” (avramidis-etal-2024-machine).
@@
For understanding computational models of language such as LLMs with billions and trillions of parameters, directed binding methods are being developed that learn to relate model internals and processes to specific human-understandable descriptions, that often lie within our vocabulary of linguistic metalanguage. We will consider a few examples of such methods that particularly use a linguistic notion of binding: rassin2023linguistic bind linguistic concepts in V&L diffusion models with syntax trees and attention maps; di-marco-etal-2023-study and jumelet-zuidema-2023-feature detect linguistic feature representations in LLMs with prompts; opitz-frank-2022-sbert bind parts of neural embedding representations to semantic aspects; geva-etal-2022-transformer view model decisions as constructed from concepts in a vocabulary. Explanations and binding can also adopt an outside, behavioral view beguvs2023large; behzad-etal-2023-elqa; chang-24. For instance, munoz2023contrasting contrast linguistic patterns in human- and LLM-generated text in morphological, syntactic, and sociolinguistic aspects. Since such methods can be also seen as diagnostic methods, the section on Evaluation (section˜4) is also relevant.
@@
On a bigger picture, linguistic metalanguage can help us understand, or establish hypotheses, for what happens in complex NLP systems, and how they differ. An example is the hypothesis that a pretrained LM can be viewed through the lens of a classical NLP pipeline, which has received arguments for and against (tenney-etal-2019-bert; de-vries-etal-2020-whats; niu-etal-2022-bert). The learning trajectory of LLMs can be traced linguistically, e.g., noting the point of syntax acquisition (chen2024sudden). LLMs have also been interpreted as “models of varieties of language” (a sociolinguistic notion, grieve2024sociolinguistic).
@@
Building a bridge to Resources (§3), we can use data sets that elicit linguistic phenomena to “open the black box” of LLMs down to the level of single neurons, as exemplified by niu2024what, who study neuron-level knowledge with the BLiMP corpus of linguistic minimal pairs (warstadt-etal-2020-blimp-benchmark). With similar means, wang-etal-2024-beyond-agreement show that BERT-like models focus on sentence-level features, whereas LLMs such as GPT or Llama are sensitive to conventions, language complexity, and organization. Experiments with complex and simple language structures can illuminate learning environments for LLMs (qin2024sometimes).
@@
Linguistically-agnostic interpretability methods can be mechanistic, i.e., formalizing the algorithms learned by neural networks pmlr-v139-weiss21a, or based on attributing a prediction to inputs via integration integrated_gradient, classifiers ribeiro2016should-lime, or Shapley values shapley1951notes. However, still, for interpreting and evaluating the results of such methods, we rely on linguistic analysis schuff-saliency-methods-need-lingustic-interpretation; feldhus-etal-2023-saliency; moeller-etal-2024-approximate, to verify that an interpretability method “outputs meaningful explanations” pmlr-v130-mardaoui21a or “emphasize[s] specific types of linguistic compositions” (kobayashi2024analyzing). This means that linguistic analyses are used to validate interpretability methods whose algorithms do not explicitly incorporate such representations.
@@
Finally, at the thousand-foot view, we see ideas from linguistics and adjacent fields take center stage in debates about how to interpret what NLP models are capable of representing—like how to define machine ‘understanding’ (dunietz-etal-2020-test; ray-choudhury-etal-2022-machine), and specifically, whether grounding is required for a model to capture meaning (e.g., bender-20; merrill-21; pavlick-23).
@@ 7 Study of Language
Our last facet highlights linguistics (and related areas) as the application domain. That is, those who study language, but are not necessarily computational linguists themselves, are a user base that can motivate NLP tasks and tools. [^10] Unlike the previous categories, which highlight ways that linguistics contributes to NLP, we view this category as bidirectional: language study motivates research towards developing NLP tools which can then contribute to the study of language. To be clear, we regard the entirety of cL—spanning formal, descriptive, distributional/statistical, and experimental inquiries—as part of linguistics. (This includes the development and use of frameworks and resources of the sorts discussed in section 3.) These inquiries contribute to linguistic understanding by advancing accounts of linguistic diversity and the precision and robustness of theories. Naturally, NLP models and representations are part of the methodology of this research (in particular, the potential relevance of LLMs to broader linguistics and cognitive science has been widely debated: see futrell-25 for a review). For purposes of this section, though, we focus on specialized NLP software that can be applied to language studies, not NLP models and representations generally. These language-system-focused applications can be distinguished from most contemporary NLP applications, which seek to make content more accessible through language (for example, by translating it, summarizing it, or reasoning about it).
@@
Classic cL tasks like parsing have applications to corpus linguistics: in order to study a linguistic pattern, it is often important to query not just strings, but with grammatical abstractions like tags, dependencies, and phrases. This was not the original goal of parsing researchers: parsing was a cL topic before the days of substantial digital corpora, and was studied to examine computational capacity to model natural language grammar, highlighting issues like formal expressive power (steedman-11). Now, the tools produced have found a secondary application in corpus linguistics, with parser outputs accessible through corpus search engines (resnik-05; fangorn; Ben:Gho:Bal:Dri:12; hundt-12; annis; grew; kulick-22). [^11] Search in parsed corpora may also be motivated by information extraction applications (shlain-20). The application to corpus linguistics motivates new angles on parsing research (e.g., syntactically searching a large text collection without having to pre-parse it in its entirety, or in a way that accounts for parser uncertainty).
@@
Documentary and historical linguistics also motivate and contribute to NLP advancement. In documentary and historical linguistics, data may be sparse, fragmented, primarily in image or audio form (without transcriptions), and lacking a standard orthography; the basic grammar of the language may still be a mystery; and the language may have few or no living speakers. (The case of endangered language documentation was discussed above in section˜5.) The exigencies of such settings call for considerably more noise-tolerance and interactivity than run-of-the-mill NLP tasks. Pure-text tasks like normalization (robertson-goldwater-2018-evaluating), as well as signal processing capabilities like OCR for historical documents (berg-kirkpatrick-etal-2013-unsupervised) or low-resource speech processing (duong-etal-2016-attentional; anastasopoulos-etal-2017-spoken; liu-etal-2022-enhancing), are important here. Algorithms and software infrastructure to assist humans engaging in multiple tasks collaboratively (elicitation, transcription, annotation, discovering grammatical generalizations, and developing usable resources for scholars and community members) in limited-data multimodal conditions are needed, and there is a long way to go before they can be used seamlessly by non-computational linguists (gessler-2022-closing; moeller-arppe-2024-machine).
@@
Language teaching, as studied in the field of Applied Linguistics and practiced in classrooms as well as other modes of instruction and technological support, is another area ripe for further engagement with NLP. Much of the research in the NLP community has been formulated in narrow tasks like grammatical error correction (as a generation task) or essay scoring (as a regression task) (burstein-99; massung-16; wang-21; klebanov-22). Applications with richer contexts—for example, technologies that would monitor student progress over time and deliver adaptive pedagogical experiences—are perhaps more difficult to study, but could enhance student learning more holistically (cui-23; qian-23; glandorf-25; bang-24).
@@
Finally, we note the potential for NLP to embrace applications facilitating language-focused study in fields beyond linguistics proper, including literature lit_language, education, law law_language, communications comm_language, translation studies translate_language, history piotrowski2012natural, argumentation cohen-1987-analyzing, and lexicography lex_language. In the NLP community, these fields are associated and developed with active workshops, e.g., LaTeCH-CLfL (CL for Cultural Heritage, Social Sciences, Humanities and Literature—8th iteration in 2024) (latechclfl-2024-joint); BEA (Building Educational Applications—19th iteration) bea-2024-innovative; Argument Mining (11th iteration) Ajjour et al. (2024); and NLLP (Natural Legal Language Processing–3rd iteration) Aletras et al. (2024), among others. Key objectives featured in these workshops include creating structured resources, increasing interpretability, evaluating models and resources, and coping with low-resource settings. This underscores the importance of interdisciplinary work between all of NLP, linguistics, and other fields from humanities and social sciences.